OpenStudy (vlery):

how would I find the exact value of TAN(-765degrees)?

OpenStudy (anonymous):

well since tan repeats itself every 360 degrees. You can just add 360 to it twice to get Tan(-45) that is Tan(-765) = Tan(-45) and tan(-x) = -tan(x) So your looking at Tan(-765) = -Tan(45) Now tan is opposite length dividied by the adjacent length (a percentage of how high you go, vs how far you go) in a triangle so when the angle in a triangle is 45, these 2 lengths are equal  and you get -Tan(45) = -a/a = -1 so Tan(-765) = -1